OpenStudy (anonymous):

What is the squared root of 1,000,000? A. 1053 B. 1032 C. 1000 D. 1019

OpenStudy (anonymous):

how many zeros you got there?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

what is half of 6?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

then your answer is a one with three zeros

OpenStudy (anonymous):

or as usual, C it is always C

OpenStudy (anonymous):

satel can you help me

OpenStudy (xapproachesinfinity):

1000000=10^6 \[\sqrt{1000000}=\sqrt{10^6}\]

OpenStudy (anonymous):

What is 5349 to the third power equals

OpenStudy (xapproachesinfinity):

see if you can do it

OpenStudy (anonymous):

lol

jimthompson5910 (jim_thompson5910):

another way to state what satellite73 is saying is to write this \[\Large 1,000,000 = 10^6\] as xapproachesinfinity wrote \[\Large \sqrt{1,000,000} = \sqrt{10^6}\] \[\Large \sqrt{1,000,000} = (10^6)^{1/2}\] \[\Large \sqrt{1,000,000} = 10^{6*1/2}\] \[\Large \sqrt{1,000,000} = 10^{6/2}\] \[\Large \sqrt{1,000,000} = 10^{3}\] \[\Large \sqrt{1,000,000} = 1,000\]

OpenStudy (anonymous):

must be ESL math "What is 5349 to the third power equals "

OpenStudy (xapproachesinfinity):

put it in calculator lol

OpenStudy (anonymous):

2.769 times 10 to the negative third power equals .002769

OpenStudy (anonymous):

.002769 cuti3

OpenStudy (anonymous):

Thanks andrew! :)

OpenStudy (anonymous):

profile picture is fake you thirsty little boys @Andrewthehelper @cuti3

OpenStudy (anonymous):

https://www.youtube.com/watch?v=ZT9egohObJU